OpenStudy (meilendurcer):

Find the surface area of the cylinder to the nearest tenth of a square unit. Use = 3.14. A. 228 cm^2 B. 399.4 cm^2 C. 165.7 cm^2 D. 371.1 cm^2

OpenStudy (meilendurcer):

@BrittanyFrench

OpenStudy (anonymous):

I know to find the volume of a cylinder you use V=BH but you are looking for the surface area, hold on one sec

OpenStudy (meilendurcer):

ok.

OpenStudy (anonymous):

okay, try 2*pi*r*h r=radius h=height

OpenStudy (anonymous):

Okay, the answer I got was not any of your choices

OpenStudy (meilendurcer):

i knew it

OpenStudy (anonymous):

? knew what?

OpenStudy (meilendurcer):

i tried and my answers weren't anywhere close

OpenStudy (anonymous):

Yeah! Mine weren't either.. trying to figure out what i did wrong

OpenStudy (anonymous):

it's B

OpenStudy (meilendurcer):

thank you

OpenStudy (anonymous):

area circle = \[\pi \times3^2=9 \pi \] \[2\times9\pi=18\pi\]area rectangle= \[6\pi \times18.2=343.06\]\[18\pi+343.06=399.61\]

OpenStudy (anonymous):

Sorry meilendurcer, I really couldn't find the answer